OpenStudy (anonymous):

A moving car decelerates for 5.0s and comes to a complete stop. It travels 75m in the process. a) determine its initial speed b)determine its rate of deceleration

OpenStudy (abmon98):

a=-am/s^2 v (final speed)=0m/s s(Distance)=75m t=5 seconds use (v+u/2)*t=s plug in the givens (u/2)*5=5 solve for u b) rate of deceleration=change in velocity/change in time=v-u/t2-t1=0-u/5